OpenStudy (anonymous):

what is the solution of the system? use substitution. 5x - y = 3 6x - y = 8 a. (1,-8) (1,-8) b. (5,22) (-5,-22) c. (11,-2) (-1,-2) d. (5/6, 3/8) (5/6, 3/8)

OpenStudy (triciaal):

to use substitution find the expression for one of the variables and use this expression in the other equation. you will have a linear equation solve for the one unknown with this value put in the expression for the other variable and compute good practice to check the solution in one of the original equations.

OpenStudy (triciaal):

for this question using eqn 1 y = 5 x - 3 put this expression for y in eqn 2 6x - y = 8 6 x - (5 x -3) = 8 x = 8 -3 = 5 y = 5(5) -3 = 22

OpenStudy (triciaal):